OpenStudy (anonymous):

Malia is using an online calculator to calculate the outputs f(n) for different inputs n. The ordered pairs below show Malia's inputs and the corresponding outputs displayed by the calculator: (1, 4), (2, 9), (3, 14), (4, 19) Which of the following functions best represents the rule that the calculator uses to display the outputs? f(n) = 5n − 1 f(n) = 5n + 1 f(n) = 4n + 1 f(n) = 4n − 1

OpenStudy (anonymous):

@madhu.mukherjee.946

OpenStudy (anonymous):

@pooja195 @ganeshie8 @Hero @Abhisar @iambatman

OpenStudy (madhu.mukherjee.946):

it is the first option

OpenStudy (madhu.mukherjee.946):

find the slope between two points do you know how to do that?

OpenStudy (madhu.mukherjee.946):

y2-y1 over x2-x1

OpenStudy (madhu.mukherjee.946):

9-4/2-1=5

OpenStudy (madhu.mukherjee.946):

now y = mx +b you got m=5 so y = 5x + b

OpenStudy (madhu.mukherjee.946):

then temporarily plug in a point (x, y) and solve for b

OpenStudy (madhu.mukherjee.946):

9=5(2) + b

OpenStudy (madhu.mukherjee.946):

-1=b

OpenStudy (madhu.mukherjee.946):

now if you consider the real equation without putting x,y it will be y=5x-1 or f(n)=5n-1 we have considered f(n)=y

OpenStudy (madhu.mukherjee.946):

y=mx+b is the equation for the straight line

OpenStudy (anonymous):

THANKSSS @madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

welcome